OpenStudy (rational):

show that \[\large \sqrt[n]{n}\lt 2\] for all \(n\gt 0\)

OpenStudy (dan815):

is there a neat solution without product log function

OpenStudy (rational):

@xapproachesinfinity claims to be having a neat solution.. he used this result in some other problem..

OpenStudy (xapproachesinfinity):

cool question, good night guys :)

OpenStudy (anonymous):

try doing it by mathematical induction

OpenStudy (sdfgsdfgs):

\[\sqrt[n]{n}<2\] as n>0, take left n right to the nth power will give n < 2^n which should be true for all n>0 am i missing something here? :)

OpenStudy (rational):

thats slick xD

OpenStudy (anonymous):

The question n^(1/n) < 2 seems algebraically equivalent to, after raising both sides to power of n, n < 2^n Now can you show that 2^n > n for all n greater than or equal to 0 ?

OpenStudy (anonymous):

@sdfgsdfgs

OpenStudy (anonymous):

it can be shown by induction i think 2^1 > 1 , basis case for n=1 Assume : 2^k > k for some k >= 1 2^(k+1) = 2 * 2^k > 2 * k > k

Nnesha (nnesha):

\(\large\color{white}{.}\)

OpenStudy (rational):

i think induction works if "n" is an integer

OpenStudy (rational):

*positive integer

OpenStudy (anonymous):

To show it is true for x continuous you can use calculus. CLaim : 2^x > x , for x > 0 :equivalent to: 2^x - x > 0 for x > 0. The derivative of the left side is positive for x > 0 , so it is increasing and remains greater than zero

OpenStudy (anonymous):

did you not mean to use positive integers n, i assumed that since you wrote n > 0

OpenStudy (rational):

positive real numbers since we're trying for rigor, how easy is it to show that \(2^x\ln(2)-1\) is always positive compared to showing that \(2^x\gt x\) ?

OpenStudy (anonymous):

2^x ln 2 -1 > 0 \( \iff \) 2^x ln 2 > 1 \( \iff \) 2^x > 1 / ln 2 \( \iff \) 2^x > 1.44 \( \iff \) if x = 1 2^1 = 2 > 1.44 and 2^x is an increasing function , which you can show by taking the derivative

OpenStudy (rational):

the derivative isn't really positive for all x>0 right ?

OpenStudy (anonymous):

2^x has a positive derivative for all x ,

OpenStudy (rational):

i thought out function is 2^x-x ?

OpenStudy (anonymous):

I wanted to demonstrate that 2^x - x is increasing function for x > 0 , and it is bounded below by 1

OpenStudy (rational):

2^x-x is increasing for sure for x > 1 but it has a local minimum in interval (0, 1)

OpenStudy (anonymous):

i want to prove that 2^x - x > 0 for all x > 0 graphically it is true for all x. but no matter 2^0 - 0 = 1 > 0

OpenStudy (anonymous):

hmm, you have a point there, it does decrease a little at first, then it increases

OpenStudy (anonymous):

ok another way to prove it is use binomial series

OpenStudy (anonymous):

2^x = (1 + 1)^x < x

OpenStudy (rational):

okay so your proof using derivatives actually works for x > 1

OpenStudy (rational):

yeah

OpenStudy (anonymous):

it works for x > c such that 2^c * ln 2 - 1 = 0

OpenStudy (rational):

yeah nice so are we using extended binomial theorem where the exponent can be a real number ?

OpenStudy (anonymous):

thats actually a fun problem, solve that equation :)

OpenStudy (anonymous):

( 1 + 1)^x = 1^x + x*1 + x(x-1)/2 * 1 + ...

OpenStudy (rational):

\[2^x=e^{x\ln 2 } = 1+x\ln 2+\frac{x^2}{2}\ln^2 2+\cdots \]

OpenStudy (rational):

just using the power series of e^x

OpenStudy (anonymous):

does that prove it is bigger than x ?

OpenStudy (rational):

hmm idk lol i don't see yet

OpenStudy (anonymous):

graphically 2^x - x > 0 for all reals , surely there must be a way to prove this analytically , for all reals

OpenStudy (rational):

yeah i like the different methods you're trying

OpenStudy (anonymous):

it is decreasing on (-oo, c) = ln(1/ln(2)) / ln (2) increasing on (c, oo)

OpenStudy (anonymous):

f(x) = 2^x -x is decreasing on (-oo, ln(1/ln(2)) / ln (2) ) increasing on ( ln(1/ln(2))/ ln (2) , oo)

OpenStudy (sdfgsdfgs):

sorry...took a shower ;) "show that 2^n > n for all n greater than or equal to 0 ? " where n is a positive real number (instead of a positive integer) since both functions, 2^n and n, are continuous, their respective values over positive real number range are bounded by the values over positive integers. Since n goes up linearly while 2^n goes up exponentially, the case is proven for all positive real numbers >1. For n between 0 and 1, 2^n will always be >1 and hence >n. So 2^n > n for all n >0 where n is a positive real no (n cannot be equal to 0 in the original inequality.)

OpenStudy (ikram002p):

I would try this Sqrt(n)(1+1+.....+1 n times ) <2

OpenStudy (anonymous):

i think its enough to show that it reaches the minimum at a positive value , and is decreasing before it, increasing after

OpenStudy (ikram002p):

So Sqrt(n)(1+1+1+1..)< n sqrt(n)(1)

OpenStudy (anonymous):

i dont know what theorem to call this approach

OpenStudy (rational):

Ahh that works! Also @sdfgsdfgs 's argument uses a simple observation : In the interval (0,1) we have 2^x > 1 and x < 1. Together imply 2^x > x. looks neat to me

OpenStudy (anonymous):

how do we prove that 2^x > 1 in (0,1)

OpenStudy (rational):

haha good q, calculus i guess

OpenStudy (sdfgsdfgs):

2^0=1, 2^1=2 n 2^n is continuous between (0,1) so 2^1 must be between 1 and 2.

OpenStudy (anonymous):

show that 2^x - 1 > 0 take derivative of 2^x -1 2^x * ln 2 -1 > 0 2^x > 1/ln 2

OpenStudy (sdfgsdfgs):

2^0=1, 2^1=2 and 2^n is continuous between (0,1) so 2^1 must be between 1 and 2.

OpenStudy (ikram002p):

Is there anyway to show that sqrt(n)(n) is not rational, I mean its clear its not integer for n>1 but testing rationality seems cool to me

OpenStudy (rational):

that looks interesting

OpenStudy (rational):

wil try a bit on the proof and post a new question :)

OpenStudy (rational):

2^0=1, 2^1=2 and 2^x is continuous between (0,1) `and 2^x increasing as its first derivative is always positive.` That means 2^x cannot go below 1 in the interval (0,1). so 2^x must be between 1 and 2.

OpenStudy (anonymous):

since the derivative of 2^x is always positive, on the interval (0,1) 2^x > 2^0 = 1> x then for x >= 1 , you can use the argument i used earlier 2^x - x > 0 take derivative of f(x) = 2^x - x this is increasing for x >=1 and f(1) = 2^1 -1 = 1 >0 therefore 2^x - x >= 1 > 0 for x >= 1 therefore 2^x > x . i hope this is complete proof

OpenStudy (rational):

looks good to me !

OpenStudy (anonymous):

maybe like this also can work: let the value of this rooth be a, then \(a^n=n\), now taking natural logs, for example, \(n\ln a= \ln n\), or \(\ln a= (\ln n)/ n\) now, getting a back: \(a=e^{(\ln n)/n}\) maybe taking ln expanssion you could figure it out from there

OpenStudy (anonymous):

i still think binomial series could prove enlightening here 2^x = (1+1)^x > x or the Bernoulli inequality

OpenStudy (anonymous):

Also can one prove from the power series expansion that $$\large 2^x=e^{x\ln 2 } = 1+x\ln 2+\frac{x^2}{2}\ln^2 2+\cdots > x $$ This is a true statement for all x

OpenStudy (xapproachesinfinity):

wow you guys really took it further:)

OpenStudy (ikram002p):

Haha yeah I was thinking its more simple like this :_ 2^n=2*...*2>=2+...+2=2n>n

OpenStudy (ikram002p):

But was having fun reading all comments that my solution sounds a bit boring :p